# UniCtrl: Improving the Spatiotemporal Consistency of Text-to-Video Diffusion Models via Training-Free Unified Attention Control

### 04 Mar 2024 v1

@@ 1 Introduction
Diffusion Models (DMs) have achieved great success in the field of image generation, demonstrating enhanced stability and superior quality compared to other generative models. Early works on DMs [46, 19, 47, 48, 31, 25] have laid the foundation for their success. These models have shown their capacity to efficiently handle and scale to diverse data, outperforming previous generative models like GANs [12, 26, 27] and VAEs [29, 55, 41] in generating images with greater stability and more impressive visual quality.
Furthermore, the evolution of DMs has introduced a significant improvement in terms of controllability and user-interactability in image generation. This progress is highlighted by the recent advancements  [42, 40, 36, 43, 65, 67, 34]. These models unleash controllability through various forms of conditions, which significantly improves user interactability, allowing for the creation of images that are more aligned with user intentions.
Recently, Video Diffusion Models (VDMs) [20] have been proposed for applying DMs to video generation tasks. With the help of text encoders [39], VDMs are capable of generating videos with a diverse range of motions in text-to-video synthesis tasks [16, 66, 22, 18, 1, 2]. Many open-source text-to-video models have been introduced, including ModelScope[58], AnimateDiff[16], VideoCrafter[6] and so on. These models typically require a pre-trained image generation model, and introduce additional temporal or motion modules. However, unlike images that contain rich semantic information, text conditions are more difficult to ensure consistency between different frames of the generated video. At the same time, some work also uses image conditions to achieve image-to-video generation with improved spatial semantic consistency [15, 23, 1]. Some works have proposed the paradigm of text-to-image-to-video[11], but image conditions alone cannot effectively control the motion of videos. Combining text and image conditions leads to enhanced spatio-temporal consistency in a text-and-image-to-video workflow [68, 6, 7, 62, 14], but these methods require additional training.
@@ 2 Related Work
Different from models that require training[67, 65, 34], attention control [17, 54, 4, 64, 9] is a training-free method which has been widely applied in the task of image editing. Previous work has found that Attention Control can be used to ensure both semantic [4] and spatial consistency [17, 54, 9] in image editing. InfEdit[64] unified the control of semantic consistency and spatial consistency for the first time, proposing unified attention control (UAC). Some work has introduced attention control to VDMs for video editing[30, 10, 28], but no one has improved the consistency of generated videos by this method. Is that possible to introduce UAC into VDMs to ensure cross-frame semantic consistency and spatial consistency throughout the video is an interesting and worthwhile question to explore.
@@ 4 UniCtrl: Cross-Frame Unified Attention Control
Figure 2: Each row contains 6 frames of videos generated with the same prompt: An Aston Martin is moving. The first row demonstrates the original frames generated with baseline model, which the vehicle and the road are clearly inconsistent across the frames. The second row shows frames generated with baseline model augmented with Cross-Frame Self-Attention Control (SAC). While it maintains semantic consistency, it exhibits little motion. The third row explains frames generated with baseline model augmented with SAC and Motion Injection (MI). Although SAC and MI ensure motion diversity and semantic consistency for each frame, they fall short on spatiotemporal consistency. The fourth row contains frames generated with baseline model and also augmented with Spatiotemporal Synchronization (SS), which improves spatiotemporal consistency over prior pipeline and achieved balance between motion and semantic, both in-frame and cross-frame.
@@
We first analyzed the role of keys and values in the self-attention layers of VDMs, and then analyzed the role of queries in all the attention layers. Inspired by InfEdit [64], we then propose the cross-frame unified attention control to achieve both semantic level consistency and better spatiotemporal consistency. Lastly, we apply spatiotemporal synchronization by replacing the latent of the motion branch with the latent from the output branch at each sampling step. We have demonstrated the effectiveness of each module of UniCtrl in Fig. 2.
@@ 4.1 Cross-Frame Self-Attention Control
Previous work [4, 17, 54, 64] has observed that queries in the attention mechanism form the layout and semantic information of generated images[4, 64, 54], while values contribute to the semantic information[17, 64]. Therefore, we hypothesize that using the same values in the attention of different frames can ensure cross-frame consistency. Additionally, we observed that the mismatch between keys and values will degrade the quality of generated videos through our experiment and we provide one qualitative example in Section 5.4. Consequently, in our cross-frame Self-Attention Control (SAC), we inject both keys and values, as detailed in Algorithm 1. We also show one example of the effectiveness of SAC in Fig. 2.
@@ 4.2 Motion Injection
Figure 3: In our framework, we use key and value from the first frame as represented by $K^{0}$ and $V^{0}$ in the self-attention block. We also use another branch to keep the motion query $Q_{\textrm{m}}$ for motion control. At the beginning of every sampling step, we let the motion latent equal to the sampling latent, to avoid spatial-temporal inconsistency.
Through our experiments in section 5, we find one key limitation of cross-frame self-attention control that may lead to overly similar frames, resulting in minimal motion in the video. Again, as described in [4, 64], the queries in self-attention and cross-attention determine the video’s spatial information, which corresponds to motion. Therefore, by preserving the original queries, we can maintain the motion effect. As Fig. 3, we have divided the inference process into two branches: the output branch, which undergoes cross-frame self-attention control, and the motion branch, which does not involve attention control. We retain the queries in the motion branch as motion queries and use the corresponding motion queries in the output branch. Here, we denote motion queries as $Q_{\textrm{m}}$. The method for motion injection can be expressed by the following formula: $out=\textsc{Attention}(Q_{\textrm{m}},\cdot,\cdot)$, where Attention refers to both of self- and cross-attention layers. We implemente two methods for motion control: 1) fully preserve the motion of the original video and 2) only preserve the motion during each step of the sampling process. To make motion more controllable, we also introduce a method to control motion through a given motion injection degree $c$ where $0\leq c\leq 1$, as following:
@@ 4.3 Spatiotemporal Synchronization
In our experiments 5, we find that the spatial information of the original video cannot ensure spatiotemporal consistency. Considering that the output of the output branch has better spatiotemporal consistency, we further propose Spatiot
emporal Synchronization (SS): the latent of the output branch is copied as the initial value of the latent for the motion branch before each sampling step. By doing so, our method can simultaneously ensure the semantic consistency of the generated video and improve the quality of spatiotemporal consistency and motion.
@@ 4.4 Cross-Frame Unified Attention Control
As shown in Fig. 3, we combine the cross-frame unified attention control with motion injection and spatiotemporal synchronization together like Unified Attention Control (UAC) in InfEdit[64]. In the output branch, we use the key $K^{0}$ and value $V^{0}$ from the first frame to ensure cross-frame semantic consistency. We use another branch to keep the reserved query for motion control, and use the query $Q_{\textrm{m}}$ from this branch to replace the query in the output branch. Before every step of sampling, we let the latent of motion branch equal to the previous output latent, to avoid spatiotemporal inconsistency. Our method can also be represented as formulated algorithm outlined in Algorithm 2.
@@ 5 Experiments
In this section, we evaluate the effectiveness of UniCtrl. The assessment includes both qualitative comparisons in Section 5.2 and quantitative comparisons in Section 5.3. Additionally, we explore the contribution of each component within UniCtrl, the motion injection degree, and the specific design choice of swapping Key and Value in the SAC procedure, as detailed in Section 5.4.
@@ 5.1 Experimental Setup
To evaluate the effectiveness of our model, we collect prompts from two datasets UCF-101[51] and MSR-VTT[63] for generating videos. Following Ge et al[8], we use the same UCF-101 prompts for our experiments. We also randomly select 100 unique prompts from the MSR-VTT dataset for our evaluation. Those two parts of data consist of our dataset for evaluations.
@@ 5.1.1 Metric
To quantitatively evaluate our results, we consider standard metrics following [44, 61]. Note that we will provide details for evaluation metrics in the supplementary material section.
@@
$\bullet$ RAFT: To compare the magnitude of actions in the videos, we utilize RAFT[52] to estimate the optical flow, thereby inferring the degree of motion. To quantify the motion intensity between every two consecutive frames, we employ the RAFT model to estimate optical flow for each pixel. We calculate the magnitude of all pixel optical flow using the $l^{2}$ norm, and finally, we obtain the average of all magnitude as the score for every consecutive frames in the video. In this process, we utilize the RAFT model from torchvision [33].
@@ 5.1.3 Baseline
We select FreeInit [61] as our baseline since FreeInit is a training free method that attempts to improve subject appearance and temporal consistency of generated videos through iteratively refining the spatial-temporal low-frequency components of the initial latent during inference. However, given that our method UniCtrl operates on the attention mechanism and FreeInit adjusts the frequency domain of the latent space, UniCtrl and FreeInit are orthogonal approaches. Both are training-free methods capable of enhancing the spatiotemporal consistency of generated videos via diffusion models. We demonstrates the integration of UniCtrl and FreeInit both in 5.2 and 5.3.
@@ 5.3 Quanlitative Comparisons
Table 1: Quantitative Comparisons on UCF-101 and MSR-VTT. UniCtrl significantly improves the temporal consistency while keeps the motion in the generated videos. $c$ indicates the motion injection degree. $I$ indicates number of iteration being used for FreeInit.
@@
For quantitative comparisons, the quantitative results on UCF-101 and MSR-VTT are reported in Table 1. We compare the backbones and backbones augmented by UniCtrl and FreeInit respectively. According to the metrics, UniCtrl significantly improves the spatiotemporal consistency in the generated videos across all backbones on both prompt sets from $0.74$ to $2.35$. While FreeInit achieves remarkable improvements over spatiotemporal consistency, we found that UniCtrl outperforms FreeInit on the diversity of motion for both AnimateDiff and VideoCrafter by a large margin from $10.91$ to $21.88$. Note that we purposely choose the motion injection degree $c=1$ to demonstrate how UniCtrl can preserve the motion compare with FreeInit. However, there still exists a trade off between spatiotemporal consistency and motion diversity. In section 5.4, we show spatiotemporal consistency can be further improved with a smaller motion injection degree. Thus, we recommend motion injection degree $c=0.4$ in real-world applications. Lastly, we showcase how UniCtrl and FreeInit can improve AnimateDiff together and we found this integration can improve both spatiotemporal consistency and motion diversity. For the details of our experiments, especially how we integrate UniCtrl and FreeInit, we will introduce the details in the supplementary material. Note that we obtained different score for FreeInit because we randomly sampled a different set of prompts from MSR-VTT. Moreover, we used our own implementations for FreeInit while conducting experiments on VideoCrafter since we cannot find official implementation.
@@ 5.4 Ablation Study
We assess the impact of each UniCtrl module, the efficacy of motion injection degree, and corroborate our design choice of swapping Key and Value during SAC through the subsequent ablation studies.
5.4.1 Evaluating the Impact of SAC, MI, and SS
To assess the contribution of the Spatial Attention Component (SAC), we conducted experiments on both datasets using AnimateDiff as the baseline, incorporating our pipeline but disabling SAC. For motion injection, we set motion injection degree to be $1$. The findings reveal our method without SAC works exactly the same the backbone because the motion branch is the exactly the same as the output branch. This observation underscores the critical role of SAC in enhancing spatiotemporal consistency, as evidenced by the comparisons with the scores obtained through the vanilla AnimateDiff.
In exploring the significance of Motion Injection (MI), additional tests were performed on both datasets with AnimateDiff serving as the baseline, this time with MI deactivated. The results indicated a notable consistency with the baseline, yet with a substantial reduction in motion diversity. This was quantitatively supported by a decrease in the RAFT score from $31.81$ to $4.19$. Such a marked disparity highlights MI’s vital contribution to maintaining motion diversity.
Table 2: Ablation on UCF-101 and MSR-VTT. We ablate each module of UniCtrl and show the effectiveness of them each. $c$ indicates the motion injection degree.
@@ 6 Conclusion
We introduce Cross-Frame Unified Attention Control (UniCtrl), to address the challenges of cross-frame consistency shift in Video Diffusion Models. By implementing UniCtrl, we’ve notably improved the alignment of semantic and spatial temporal consistency across frames of generated videos. Our approach stands out as it requires no additional training, making it adaptable to various underlying models. The efficacy of UniCtrl has been rigorously tested, demonstrating its potential to be widely applied to text-to-video generative models. We introduce some key limitations of UniCtrl in section 6.1 and our ethics statement in section 6.2.
@@ 6.2 Ethics Statement
7 Acknowledgment
We thank Raven, Tom, Martin, Yichi Zhang, Shengyi Qian for their helpful feedback and advice.

### 06 Mar 2024 v3

@@ 1 Introduction
Diffusion Models (DMs) have excelled in image generation, offering enhanced stability and quality over methods like GANs [12, 26, 27] and VAEs [30, 57, 42]. Foundational studies [48, 19, 49, 50, 32, 25] established the groundwork for DMs’ capabilities in efficiently scaling up with diverse data. Recent advancements [43, 41, 37, 45, 67, 69, 35] have further improved controllability and user interaction, enabling the creation of images that better reflect user intentions.
Recently, Video Diffusion Models (VDMs) [20] have been proposed for utilizing DMs for video generation tasks. VDMs are capable of generating videos with a wide variety of motions in text-to-video synthesis tasks, supported by the integration of text encoders [40], as shown in [16, 68, 22, 18, 1, 2]. Many open-source text-to-video models have been introduced, including ModelScope[60], AnimateDiff[16], VideoCrafter[6] and so on. These models typically require a pre-trained image generation model, e.g., Stable Diffusion(SD) [44], and introduce additional temporal or motion modules. However, unlike images that contain rich semantic information, text conditions are more difficult to ensure consistency between different frames of the generated video. At the same time, some work also uses image conditions to achieve image-to-video generation with improved spatial semantic consistency [15, 23, 1]. Some works have proposed the paradigm of text-to-image-to-video[11], but image conditions alone cannot effectively control the motion of videos. Combining text and image conditions leads to enhanced spatiotemporal consistency in a text-and-image-to-video workflow [70, 6, 7, 64, 14], but these methods require additional training.
@@ 2 Related Work
Different from models that require training[69, 67, 35], attention control [17, 56, 4, 66, 9] is a training-free method which has been widely applied in the task of image editing. Previous work has found that Attention Control can be used to ensure both semantic [4] and spatial consistency [17, 56, 9] in image editing. InfEdit[66] unified the control of semantic consistency and spatial consistency for the first time, proposing unified attention control (UAC). Text2Video-Zero [28] applies frame-level self-attention on text-to-image synthesis methods and enables text-to-video through manipulating motion dynamics in latent codes. Some work has introduced attention control to VDMs for video editing[31, 10, 29], but no one has improved the consistency of generated videos through video diffusion by attention control. Is that possible to introduce UAC into VDMs to ensure cross-frame semantic consistency and spatial consistency throughout the video is an interesting and worthwhile question to explore.
@@ 4 UniCtrl: Cross-Frame Unified Attention Control
Figure 2: The first row demonstrates the original frames generated with baseline model, which the vehicle and the road are clearly inconsistent across the frames. The second row shows frames generated with baseline model augmented with Cross-Frame Self-Attention Control (SAC). While it maintains incredible spatiotemporary consistency, it exhibits little motion. The third row explains frames augmented with SAC and Motion Injection (MI). Although MI injects more motion in addition to SAC, the results clear demonstates that it falls short on spatiotemporal consistency again. The fourth row contains frames further augmented with Spatiotemporal Synchronization (SS) in addition to SAC and MI, which improves spatiotemporal consistency over the results from the thrid row and achieved balance between motion and spatiotemporary consistency, both in-frame and cross-frame.
@@
We first analyzed the role of keys and values in the self-attention layers of VDMs, and then analyzed the role of queries in all the attention layers. Inspired by InfEdit [66], we then propose the cross-frame unified attention control to achieve both semantic level consistency and better spatiotemporal consistency. Lastly, we apply spatiotemporal synchronization by replacing the latent of the motion branch with the latent from the output branch at each sampling step. We demonstate the effectiveness of each module of UniCtrl in Fig. 2.
@@ 4.1 Cross-Frame Self-Attention Control
Previous work [4, 17, 56, 66] has observed that queries in the attention mechanism form the layout and semantic information of generated images[4, 66, 56], while values contribute to the semantic information[17, 66]. Therefore, we hypothesize that using the same values in the attention of different frames can ensure cross-frame consistency. Additionally, we observed that the mismatch between keys and values will degrade the quality of generated videos through our experiment and we provide one qualitative example in Section 5.4. Consequently, in our cross-frame Self-Attention Control (SAC), we inject both keys and values from the first frame, as detailed in Algorithm 1, to every other frame at each self-attention layer during denoising. We showcase one example of the effectiveness of SAC by comparing the first row and the second row in Fig. 2.
@@ 4.2 Motion Injection
Figure 3: In our framework, we use key and value from the first frame as represented by $K^{0}$ and $V^{0}$ in the self-attention block. We also use another branch to keep the motion query $Q_{\textrm{m}}$ for motion control. At the beginning of every sampling step, we let the motion latent equal to the sampling latent, to avoid spatial-temporal inconsistency. We explain details of our framework in Algorithm 1 and Algorithm 2.
In our experiments, detailed in Section 5, we identify a significant limitation associated with cross-frame self-attention control: it tends to produce overly similar consecutive frames, leading to minimal motion within the video sequence. Again, as described in [4, 66], the queries in self-attention and cross-attention determine the video’s spatial information, which corresponds to motion. Therefore, by preserving the original queries, we can maintain the motion effect.
As in Fig. 3, we have divided the inference process into two branches: the output branch, which undergoes cross-frame self-attention control, and the motion branch, which does not involve attention control. We retain the queries in the motion branch as motion queries and use the corresponding motion queries in the output branch. Here, we denote motion queries as $Q_{\textrm{m}}$. The method for motion injection can be expressed by the following formula: $out=\textsc{Attention}(Q_{\textrm{m}},\cdot,\cdot)$, where Attention refers to both of self- and cross-attention layers. This method is designed to fully retain the motion of the original video. However, in practice, a trade-off between motion preservation and spatiotemporal coherence remains evident. Consequently, we propose an additional refinement: selectively preserving motion at specific steps throughout the sampling process to enhance control. To this end, we introduce a technique that modulates motion through the integration of a motion injection degree, $c$, which is defined within the interval $0\leq c\leq 1$. The following outlines this approach in detail:
@@ 4.3 Spatiotemporal Synchronization
Upon closer inspection of the results, notably exemplified by the third row of Figure 2, it becomes evident that simply injecting spatial information from the original video is insufficient for guaranteeing spatiotemporal consistency. Considering that output branch yields more spatiotemporally consistent results, we further propose Spatiotemporal Synchronization (SS): the latent of the output branch is copied as the initial value of the latent for the motion branch before each sampling step. By doing so, our method can simultaneously ensure the semantic consistency of the generated video and improve the quality of spatiotemporal consistency and preserve the degree of motion diversity. We present a qualitative example to demonstrate the effectiveness of SS by comparing the results depicted in the third and fourth rows of Figure 2.
@@ 4.4 Cross-Frame Unified Attention Control
As illustrated in Figure 3, we integrate Cross-Frame Self-Attention Control, Motion Injection and Spatiotemporal Synchronization into a cohesive framework termed Cross-Frame Unified Attention Control. In the output branch, we employ the key $K^{0}$ and value $V^{0}$ derived from the initial frame to maintain cross-frame semantic consistency. A separate branch is utilized to preserve the query specifically for motion control, replacing the output branch’s query with $Q_{\text{m}}$ from the motion control branch. To prevent spatiotemporal inconsistencies, we synchronize the latent representation of the motion branch with the output branch’s preceding latent state before each sampling step. This approach is further detailed and formalized presented in Algorithm 2.
@@ 5 Experiments
In this section, we evaluate the effectiveness of UniCtrl. We discuss metrics, backbones and baseline in Section 5.1. Then we include both qualitative comparisons in Section 5.2 and quantitative comparisons in Section 5.3 to showcase the effectiveness of UniCtrl in terms spatiotemporal consistency and motion preservation. Additionally, we explore the contribution of each component within UniCtrl, the motion injection degree, and the specific design choice of swapping Key and Value together in the SAC procedure, as detailed in Section 5.4.
@@ 5.1 Experimental Setup
To evaluate the effectiveness of our model, we collect prompts from two datasets UCF-101[53] and MSR-VTT[65] for generating videos. Following Ge et al[8], we use the same UCF-101 prompts for our experiments. We also randomly select 100 unique prompts from the MSR-VTT dataset for our evaluation. Those two parts of data consist of our dataset for evaluations. Next we briefly introduce evaluation metrics and we will provide details in the supplementary material.
@@ 5.1.1 Metric
To quantitatively evaluate our results, we consider standard metrics following [46, 63]:
@@
$\bullet$ RAFT: To compare the magnitude of motion in the videos, we utilize RAFT[54] to estimate the optical flow, thereby inferring the degree of motion. We utilize the off-the-shelf RAFT model from torchvision [34].
@@ 5.1.3 Baseline
We select FreeInit [63] as our baseline since FreeInit is a training free method that attempts to improve subject appearance and temporal consistency of generated videos through iteratively refining the spatial-temporal low-frequency components of the initial latent during inference. However, given that our method UniCtrl operates on the attention mechanism and FreeInit adjusts the frequency domain of the latent space, UniCtrl and FreeInit are orthogonal approaches. Both are training-free methods capable of enhancing the spatiotemporal consistency of generated videos via diffusion models. We demonstrate the integration of UniCtrl and FreeInit both in 5.2 and 5.3.
@@ 5.3 Quanlitative Comparisons
Table 1: Quantitative Comparisons on UCF-101 and MSR-VTT. UniCtrl significantly improves the temporal consistency while keeps the motion in the generated videos. $c$ indicates the motion injection degree. $I$ indicates number of iteration for FreeInit.
@@
For quantitative comparisons, the quantitative results on UCF-101 and MSR-VTT are reported in Table 1. We compare the backbones and backbones augmented by UniCtrl and FreeInit respectively. According to the metrics, UniCtrl significantly improves the spatiotemporal consistency in the generated videos across all backbones on both prompt sets from $0.74$ to $2.35$. While FreeInit achieves remarkable improvements over spatiotemporal consistency, we found that UniCtrl outperforms FreeInit on the strength of motions on both AnimateDiff and VideoCrafter by a large margin from $10.91$ to $21.88$. Note that we purposely choose the motion injection degree $c=1$ to demonstrate how UniCtrl can preserve the motion compare with FreeInit. However, there still exists a trade off between spatiotemporal consistency and motion diversity. In section 5.4, we show spatiotemporal consistency can be further improved with a smaller motion injection degree. Thus, we recommend motion injection degree $c=0.4$ for real-world applications. Lastly, we showcase how UniCtrl and FreeInit can improve AnimateDiff together and we found this integration can improve both spatiotemporal consistency and motion diversity. We will introduce the details how we integrate UniCtrl and FreeInit in the supplementary material. Note that we obtained different score for FreeInit because we randomly sampled a different set of prompts from MSR-VTT and we used our own implementations for FreeInit on VideoCrafter since we cannot find official implementation.
@@ 5.4 Ablation Study
Table 2: Ablation results on UCF-101 and MSR-VTT. We ablate each module of UniCtrl and show the effectiveness of them each. $c$ indicates the motion injection degree.
@@
We assess the impact of each UniCtrl module, the efficacy of motion injection degree, and corroborate our design choice of swapping Key and Value together in the SAC procedure through the subsequent ablation studies.
5.4.1 Evaluating the Impact of SAC, MI, and SS
To assess the contribution of the SAC, we conducted experiments on both datasets using AnimateDiff as the baseline, incorporating our pipeline but disabling SAC. For motion injection, we set motion injection degree to be $1$. The findings reveal our method without SAC works exactly the same the backbone because the motion branch is the exactly the same as the output branch. This observation underscores the critical role of SAC in enhancing spatiotemporal consistency, as evidenced by the comparisons with the scores fom the vanilla AnimateDiff.
In exploring the significance of Motion Injection (MI), additional tests were performed on both datasets with AnimateDiff serving as the baseline, this time with MI deactivated. The results indicated a notable consistency with the baseline, yet with a substantial reduction in motion diversity. This was quantitatively supported by a decrease in the RAFT score from $31.81$ to $4.19$. Such a marked disparity highlights MI’s vital contribution to maintaining the motion dynamics.
@@ 6 Conclusion
We introduce UniCtrl to address the challenges of maintaining cross-frame consistency and preserving motions for Video Diffusion Models. By incorporating UniCtrl, we have notably improved the spatiotemporal consistency across frames of generated videos. Our approach stands out as it requires no additional training, making it adaptable to various underlying models. The efficacy of UniCtrl has been rigorously tested, demonstrating its potential to be widely applied to text-to-video generative models. We discuss the primary limitations of UniCtrl in Section 6.1 and detail our ethics statement in Section 6.2.